Given quantity demanded t



And supplied as q=36-1/3p and q=- 9+1/2p. If the government decide to impose a tax of 't" per unit supplied determine the value of tax

1
Expert's answer
2022-03-14T10:30:27-0400

"q=36-\frac{1}{3}p" (No tax)

"q=36-\frac{1}{3}(p-t)" with tax

"q=- 9+\frac{1}{2}p"

"Q_s= Q_d"

"36-\frac{1}{3}(p-t)= - 9+\frac{1}{2}p"

"45+ \frac{1}{3}t= \frac{1}{3}P+ \frac{1}{2}P"

"\frac{1}{3}t= \frac{5}{6}P- 45"

"P= (45+ \frac{1}{3}t)\frac {6}{5}"

"P= 80+ \frac{2}{5}t" .... price paid by consumers

"\therefore" Consumers pay "\frac{2}{5}" and the suppliers "\frac{3}{5}"